If 40mg of a radioactive substance dcays to 5mg in 12min, find the half-life, in minutes of the substance.
The halflife formula is
N = No (1/2) ^t/h
so i think that the formula goes to
5 = 40 (1/2) ^ 5/h but im not sure if thats right and i can't figure out how to solve

Answer by stanbon(48569)

If 40mg of a radioactive substance dcays to 5mg in 12min, find the half-life, in minutes of the substance.
------------
Decay formula: A(t) = Ao(x)^t
5 = 40(x)^12
1/8 =x^12
x=0.841
--------
EQUATION:
A(t) = Ao (0.841)^t
--------
Now find the half-life:
(1/2)Ao = Ao(0.841)^t
(1/2)=(0.841)^t
t=ln(1/2) / ln0.841 = 4 minutes
----------
The half-life is 4 minutes.
==========
Cheers,
Stan H.